OpenStudy (anonymous):

the quotient given below arises in calcullus and is not defined when h=0. Verify the identity given below: [sin(x+h)-sin(x)]/h = cos(x(sin(h)/h)) - sin(x(1-cos(h)/h)))

OpenStudy (anonymous):

\[(\sin(x+h) - \sin(x)) / h = \cos x ((\sin h)/h) - \sin x ((1-\cos h) /h)\]